OpenStudy (anonymous):

give the slope-intercept form of equation of the line that is perpendicular to -2x+3y=10 and contains P 10,-10

OpenStudy (shamil98):

Perpendicular lines have inverse slopes. First, turn that equation into slope-intercept form.

OpenStudy (shamil98):

y = mx + b is slope intercept form..

OpenStudy (anonymous):

Change -2x +3y = 10 into the slope intercept form, which is y=mx + b. So it would be y= 2/3 x + 10/3. The slope will the the opposite reciprocal, so the slope would be -3/2 x. Plug in your points, (10, -10) for x and y. So it would be: -10=-3/2(10) + b Simplify and you get b=5. So the answer will be y=-3/2x + 5